OpenStudy (anonymous):

Find the cube of 20. A. 40 B. 400 C. 800 D. 8000

OpenStudy (anonymous):

\[x^3 = x \times x \times x\] Now if \(x\) is \(20\), then what is \((20)^3\) ??

OpenStudy (anonymous):

\(x^3\) is called cube of \(x\)..

OpenStudy (anonymous):

@cinaeth your keyboard is alright no?

OpenStudy (anonymous):

it's D; 8000